Peter bought a new graduated cylinder for his chemistry class. It holds 490 mL of liquid. If the cylinder has a radius of 3.5 cm, then how tall is the cylinder?


1
Expert's answer
2021-05-07T11:42:14-0400

"r=3.5cm,"


Volume of cylinder"=490ml=490cm^3"


"\Rightarrow \pi r^2h=490"


"\Rightarrow h=\dfrac{490}{(3.5)^2\times 3.14}=12.73cm"


Hence The cylinder must be 12.73 cm tall to hold 490 ml of liquid.